# Bosenovae with Quadratically-Coupled Scalars in Quantum Sensing Experiments

### v1

@@ Abstract
Ultralight dark matter (ULDM) particles of mass $m_{\phi}\lesssim 1~{}{\rm eV}$ can form boson stars in DM halos. Collapse of boson stars leads to explosive bosenova emission of copious relativistic ULDM particles. In this work, we analyze sensitivity of terrestrial and space-based experiments to detect such relativistic scalar ULDM particles interacting through quadratic couplings with Standard Model constituents, including electrons, photons and gluons. We highlight key differences with searches for linear ULDM couplings. Screening of ULDM with quadratic couplings near the surface of the Earth can significantly impact observations in terrestrial experiments, motivating future space-based experiments. We demonstrate excellent ULDM discovery prospects, especially for quantum sensors, which can probe quadratic couplings orders below existing constraints by detecting bosenova events in the ULDM mass range $10^{-23}\,{\rm eV}\lesssim m_{\phi}\lesssim 10^{-5}\,{\rm eV}$. We also report updated constraints on quadratic couplings of ULDM in case it comprises cold DM.
preprint: KEK-QUP-2024-0002, KEK-TH-2599, KEK-Cosmo-0338, IPMU24-0004
I Introduction
@@
The predominant constituent of matter in the Universe, contributing $\sim$85%, resides in mysterious dark matter (DM) (see e.g. Ref. Cooley:2022ufh for an overview). Despite decades of efforts, DM continues to evade direct detection aiming to detect DM energy depositions in experiments due to non-gravitational DM interactions. Significant attention has been devoted to exploring weakly interacting massive particles, which have typical masses in the range of $\sim 1$ GeV – 100 TeV (see e.g. Ref. Roszkowski:2017nbc ; Schumann:2019eaa for reviews). A well-motivated class of DM that has been less studied is ultralight dark matter (ULDM) ULDM , having mass of $m_{\phi}\lesssim 1\,{\rm eV}$. With such small constituent masses, the DM is characterised by very large occupation numbers and behaves like a coherent classical wave (e.g. Hu:2000ke ; Hui:2016ltb ). The associated direct DM detection signatures are also distinct.
@@
In the case of scalar ULDM, direct couplings with the Standard Model (SM) can lead to the variation of fundamental constants, such as the fine-structure constant $\alpha$. The experimental program searching for scalar ULDM has particularly benefited from variety of approaches based on high-precision instruments ULDM ; Antypas:2022asj ; adams2023axion , such as atomic clocks Beloy:2020tgz ; Filzinger:2023zrs ; Sherrill:2023zah ; 2022PhRvL.129x1301K ; Hees:2016gop ; Arvanitaki:2014faa ; Aharony:2019iad ; Banerjee:2023bjc ; SrOH2021 ; Zhang:2022ewz ; Zaheer:2023ulf , optical cavities Kennedy:2020bac ; Tretiak:2022ndx ; Campbell:2020fvq ; Geraci:2018fax , spectroscopy Oswald:2021vtc ; VanTilburg:2015oza , mechanical resonators Branca:2016rez ; Arvanitaki:2015iga ; Manley:2019vxy , optical interferometers Savalle:2020vgz ; Aiello:2021wlp ; Vermeulen:2021epa ; Fukusumi:2023kqd , atom interferometers Badurina:2021rgt ; 2021QS&T….6d4003A , tests of the equivalence principle (EP) Hees:2018fpg ; Berge:2017ovy , and fifth force searches Fischbach:1996eq ; Adelberger:2003zx ; Murata:2014nra whose complementary sensitivities span and cover a wide ULDM mass range.
@@
ULDM can form gravitationally-bound boson stars Kaup:1968zz ; Ruffini:1969qy ; Colpi:1986ye through both gravitational interactions Levkov:2018kau ; Chen:2020cef and the self-interactions of the ULDM field Kirkpatrick:2020fwd ; Chen:2021oot ; Kirkpatrick:2021wwz . In the presence of attractive self-interactions, the boson star is expected to become unstable once it reaches a critical mass. This occurs when the density of the boson star increases enough for the ULDM self-interactions to destablize the previously-established hydrostatic equilibrium between the boson star’s gravitational and the gradient energy Chavanis:2011zi ; Chavanis:2011zm ; Eby:2014fya . Subsequently, the boson star collapses and, due to the same self-interactions, explodes in a bosenova event, copiously emitting relativistic ULDM scalars Eby:2016cnq ; Levkov:2016rkk ; Helfer:2016ljl ; Eby:2017xrr . Due to wave-spreading of relativistic scalar bursts in-flight away from the source the signal can persist for a significant time, even months to years, as it traverses terrestrial or space-based experiments Eby:2021ece . In previous work Arakawa:2023gyq , we explored transient signals from bosenovae, establishing prospects for enhanced reach in case of dilatonic ULDM couplings to SM present at first order in the effective field theory (EFT) description. Some of us also have explored prospects for detection of bosenovae from axion stars Eby:2021ece . Complementary, relativistic particles originating from historic bosenova events can contribute to diffuse background with distinct observable signatures Eby:2024mhd .
@@
In this work we explore the detection prospects for bosenovae focusing on quadratic couplings of ULDM to the SM. Quadratic couplings could dominate over linear ULDM couplings in scenarios with additional symmetries, such as a $\mathbb{Z}_{2}$ symmetry under which ULDM $\phi$ field is odd. This is readily realized for axions, which transform as a pseudoscalar field under parity symmetry. As a consequence, axion-like fields including the QCD axion generically possess quadratic couplings without corresponding linear “dilatonic” coupling Kim:2022ype ; Beadle:2023flm ; Kim:2023pvt . In contrast to the analysis of observational bosenova signatures with linear ULDM couplings to SM Arakawa:2023gyq , quadratic interactions posses additional phenomenological consequences such as screening of the field in the presence of matter density Hees:2018fpg .
@@
The paper is organized as follows. First, in Sec. II we describe the EFT characterizing the ULDM-SM interactions, as well as the self-interaction coupling. Then, in Sec. III we give an overview of boson stars and bosenovae which provide the astrophysical signal. Next, in Sec. IV we give an outline of the detection methods for bosenovae and experimental sensitivities. In Sec. V we describe implications of our results. We conclude in Sec. VI.
II Theoretical Model
II.1 Effective interactions
@@
where $m_{\phi}$ is the field’s mass and $\lambda$ is the self-interaction coupling, and the []delimited-[][$\dots$][ … ] indicate possible terms with higher powers of $\phi$. Here, we consider $\phi$ to constitute ULDM with mass $m_{\phi}\lesssim{\rm eV}$, and treat $|\lambda|\ll 1$ as a free parameter that depends on the underlying theory. We focus on parameter space $\lambda>0$, denoting attractive self-interactions between $\phi$ particles [^1] This is also motivated by some theoretical considerations, see e.g. Fan:2016rda .. In Sec. II.3 we discuss various constraints on $\lambda$.
Couplings of $\phi$ to the SM, and as a consequence their optimal methods for detection, will depend on the symmetries such as parity of the model. Axions and axion-like particles more generally (see e.g. adams2023axion ) are typically parity-odd, which implies derivative couplings to SM fields. On the other hand, parity-even fields (see e.g. Antypas:2022asj ) have different interactions with SM fields. As we discuss below, this leads to distinct detection prospects and promising search strategies.
@@
where $M_{\text{pl}}=1.2\times 10^{19}\,{\rm GeV}$ is the Planck mass and ${\mathcal{O}}_{\rm SM}^{i}$ are dimension-four SM operators labeled by index $i$. In this parametrization the $d_{i}^{(n)}$ couplings are dimensionless. Throughout, we will focus on scalars with quadratic $\propto\phi^{2}$ interactions with SM constituents, corresponding to Eq. (2) with $n=2$. In particular we consider
$\mathcal{L}_{\text{int}}\supset$ $~{}\frac{4\pi\,\phi^{2}}{M_{\text{pl}}^{2}}\bigg{(}d_{m_{e}}^{(2)}m_{e}\overline{e}e-\frac{d_{e}^{(2)}}{4}F_{\mu\nu}F^{\mu\nu}$
$-\frac{d_{g}^{(2)}\beta(g_{s})}{2g_{s}}G_{\mu\nu}^{a}G^{a\mu\nu}\bigg{)}~{},$ (3)
@@
In the presence of the couplings described in Eq. (II.1), classical oscillations of the ULDM field at frequency of order $m_{\phi}$ lead to oscillations of the fundamental constants, $m_{e}/m_{p}$ (electron-to-proton mass ratio), $\alpha$ (fine structure constant), and $m_{q}/\Lambda_{\rm QCD}$ (ratio of quark mass to QCD scale), respectively. As we discuss, this allows for unique opportunities to search for their manifestations in quantum sensing experiments. Typically, such searches focus on detecting local DM abundance in the vicinity of Earth, which is expected to have an energy density of order $\rho_{\rm DM}\simeq 0.4\,{\rm GeV/cm}^{3}$ Read:2014qva . However, burst emissions of relativistic ULDM arising from Galactic transient sources establish novel and unique complementary observational targets Eby:2021ece ; Arakawa:2023gyq . Previously Arakawa:2023gyq , we analyzed the detection prospects for ULDM with linear $\propto\phi$ interactions with the SM constituents, corresponding to $n=1$ case in Eq. (2). In this study, we analyze the signatures and detection prospects for ULDM scalars with quadratic SM couplings.
II.2 Screening
@@
II.3 Constraints on self-interaction coupling $\lambda$
II.3.1 Observational constraints
@@ II.3.2 Naturalness
Figure 2: One-loop diagrams that generate corrections to ULDM self-interaction coupling $\lambda$ originating from the quadratic couplings to SM $d_{m_{e}}^{(2)}$ (top left), $d_{e}^{(2)}$ (top right) and $d_{g}^{(2)}$ (bottom). See Eq. (II.1) of text.
@@
$\lambda(\Lambda_{\text{UV}})\simeq$ $~{}\lambda(\mu_{0})-\frac{(d^{(2)}_{m_{e}})^{2}m_{e}^{2}}{4\pi^{2}M_{\text{pl}}^{4}}\Lambda_{\text{UV}}^{2}$
$+\frac{(d^{(2)}_{e})^{2}}{8\pi^{2}M_{\text{pl}}^{4}}\Lambda_{\text{UV}}^{4}+\frac{2(d^{(2)}_{g}\beta(g_{s}))^{2}}{\pi^{2}g_{s}^{2}M_{\text{pl}}^{4}}\Lambda_{\text{UV}}^{4}\,,$ (13)
where $\mu_{0}\ll\Lambda_{\rm UV}$ is a reference energy scale. The same couplings in Eq. (II.1) imply divergent contributions $\delta m_{\phi}^{2}$ to the the ULDM mass $m_{\phi}$ (see e.g. Craig:2022eqo for a recent review). Requiring a “natural” ULDM mass, with corrections being subdominant $\delta m_{\phi}\lesssim m_{\phi}$, implies
@@
Applying these limits to Eq. (II.3.2), the naturalness limits on renormalized self-couplings are
@@
Figure 3: Constraints on ULDM self-interaction $\lambda$ versus mass $m_{\phi}$. The bounds from the Bullet Cluster (cyan) Markevitch:2003at , structure formation (orange) Cembranos:2018ulm , Lyman-$\alpha$ Irsic:2017yje ; Rogers:2020ltq and ultra-faint dwarf (UFD) galaxies Dalal:2022rmp (red vertical thick and dashed lines, respectively), black hole superradiance Baryakhtar:2020gao ; Unal:2020jiy (BHSR, brown), as well as the lower bounds for naturalness arising from $d^{(2)}_{m_{e},e}$ (red), and $d^{(2)}_{g}$ (pink) are shown. The naturalness lines assume benchmark values for the couplings that are allowed by current experimental observations, as labeled. We also illustrate the region where critical boson stars would not undergo bosenova due to general relativistic effects (purple hatched region). The dashed black line represents the benchmark $\lambda_{b}$ that we choose at each mass, as given in Eq. (12) and discussed in the Sec. II.3. The dotted black line shows the benchmark $\lambda_{b}^{\prime}$ we choose when we consider the BHSR bounds, given in Eq. (30). See text for details.
The constraints from naturalness are shown in Fig. 3 for a benchmark choice of the dilatonic coupling $d_{i}^{(2)}=10^{-4}$, which is allowed by all current experimental bounds. We note that dilatonic couplings that are larger than this are also allowed by experimental constraints at the higher end of the mass spectrum (see Sec. V for details), and also the dilatonic couplings could be significantly smaller. The naturalness constraint depends on the choice of dilatonic coupling linearly, as can be seen from Eqs. (17-19). While we illustrate constraints from naturalness considerations, we do not consider them here as strict limits for our choice of benchmark parameter values of self-interaction coupling $\lambda$.
III Boson Stars and Bosenovae

### v2

@@ Abstract
Ultralight dark matter (ULDM) particles of mass $m_{\phi}\lesssim 1~{}{\rm eV}$ can form boson stars in DM halos. Collapse of boson stars leads to explosive bosenova emission of copious relativistic ULDM particles. In this work, we analyze the sensitivity of terrestrial and space-based experiments to detect such relativistic scalar ULDM particles interacting through quadratic couplings with Standard Model constituents, including electrons, photons, and gluons. We highlight key differences with searches for linear ULDM couplings. Screening of ULDM with quadratic couplings near the surface of the Earth can significantly impact observations in terrestrial experiments, motivating future space-based experiments. We demonstrate excellent ULDM discovery prospects, especially for quantum sensors, which can probe quadratic couplings orders below existing constraints by detecting bosenova events in the ULDM mass range $10^{-23}\,{\rm eV}\lesssim m_{\phi}\lesssim 10^{-5}\,{\rm eV}$. We also report updated constraints on quadratic couplings of ULDM in case it comprises cold DM.
arxiv: 2402.06736preprint: KEK-QUP-2024-0002, KEK-TH-2599, KEK-Cosmo-0338, IPMU24-0004
1 Introduction
@@
The predominant component of matter in the Universe, contributing $\sim$85%, resides in mysterious dark matter (DM) (see e.g. Ref. Cooley:2022ufh for an overview). Despite decades of effort, DM continues to evade direct detection aiming to detect DM energy depositions in experiments due to non-gravitational DM interactions. Significant attention has been devoted to exploring weakly interacting massive particles (WIMPs), which have typical masses in the range of $\sim 1$ GeV – 100 TeV (see e.g. Ref. Roszkowski:2017nbc ; Schumann:2019eaa for reviews). A well-motivated class of DM that has been less studied is ultralight dark matter (ULDM) ULDM , having mass $m_{\phi}\lesssim 1\,{\rm eV}$. With such small $m_{\phi}$, the DM is characterised by very large occupation numbers and behaves like a coherent classical wave (e.g. Hu:2000ke ; Hui:2016ltb ). The associated direct DM detection signatures are also distinct.
@@
In the case of scalar ULDM, direct couplings with the Standard Model (SM) can lead to the variation of fundamental constants, such as the fine-structure constant $\alpha$ Arvanitaki:2014faa ; PhysRevLett.115.201301 . The experimental program searching for scalar ULDM has particularly benefited from variety of approaches based on high-precision instruments ULDM ; Antypas:2022asj ; adams2023axion , such as atomic clocks PhysRevLett.117.271601 ; Beloy:2020tgz ; Filzinger:2023zrs ; Sherrill:2023zah ; 2022PhRvL.129x1301K ; Hees:2016gop ; Arvanitaki:2014faa ; Aharony:2019iad ; Banerjee:2023bjc ; SrOH2021 ; Zhang:2022ewz ; Zaheer:2023ulf , optical cavities Kennedy:2020bac ; Tretiak:2022ndx ; Campbell:2020fvq ; Geraci:2018fax , spectroscopy Oswald:2021vtc ; VanTilburg:2015oza , mechanical resonators Branca:2016rez ; Arvanitaki:2015iga ; Manley:2019vxy , optical interferometers PhysRevLett.114.161301 ; PhysRevA.93.063630 ; Grote:2019uvn ; Aiello:2021wlp ; Vermeulen:2021epa ; Fukusumi:2023kqd , atom interferometers Badurina:2021rgt ; 2021QS&T….6d4003A , tests of the equivalence principle (EP) Hees:2018fpg ; Berge:2017ovy , and fifth force searches Fischbach:1996eq ; Adelberger:2003zx ; Murata:2014nra whose complementary sensitivities span and cover a wide ULDM mass range.
@@
ULDM can form gravitationally-bound boson stars Kaup:1968zz ; Ruffini:1969qy ; Colpi:1986ye through both gravitational interactions Levkov:2018kau ; Chen:2020cef and the self-interactions of the ULDM field Kirkpatrick:2020fwd ; Chen:2021oot ; Kirkpatrick:2021wwz . In the presence of attractive self-interactions, the boson star is expected to become unstable once it reaches a critical mass. This occurs when the density of the boson star increases enough for the ULDM self-interactions to destablize the previously-established hydrostatic equilibrium between the boson star’s gravitational and the gradient energy Chavanis:2011zi ; Chavanis:2011zm ; Eby:2014fya . Subsequently, the boson star collapses and, due to the same self-interactions, explodes in a bosenova event, copiously emitting relativistic ULDM scalars Eby:2016cnq ; Levkov:2016rkk ; Helfer:2016ljl ; Eby:2017xrr . Due to wave-spreading of relativistic scalar bursts in-flight away from the source the signal can persist for a significant time, even months to years, as it traverses terrestrial or space-based experiments Eby:2021ece . In previous work Arakawa:2023gyq , we explored transient signals from bosenovae, establishing prospects for enhanced reach in case of dilatonic ULDM couplings to SM present at first order in the effective field theory (EFT). Some of us also have explored the prospects for the detection of bosenovae from axion stars Eby:2021ece . Complementarily, relativistic particles originating from historic bosenova events can contribute to diffuse background with distinct observable signatures Eby:2024mhd .
@@
In this work we explore the detection prospects for bosenovae focusing on quadratic couplings of ULDM to the SM. Quadratic couplings could dominate over linear ULDM couplings in scenarios with additional symmetries, such as a $\mathbb{Z}_{2}$ symmetry under which ULDM $\phi$ field is odd. This is readily realized for axions, which transform as a pseudoscalar field under parity symmetry. As a consequence, axion-like fields including the QCD axion generically possess quadratic couplings without corresponding linear “dilatonic” coupling Kim:2022ype ; Beadle:2023flm ; Kim:2023pvt . In contrast to the analysis of observational bosenova signatures with linear ULDM couplings to SM Arakawa:2023gyq , quadratic interactions posses additional phenomenological consequences, such as screening of the field in the presence of ordinary matter Hees:2018fpg .
@@
The paper is organized as follows. First, in Sec. 2 we describe the EFT characterizing the ULDM-SM interactions, as well as the self-interaction coupling. Then, in Sec. 3 we give an overview of boson stars and bosenovae which provide the astrophysical signal. Next, in Sec. 4 we give an outline of the detection methods for bosenovae and experimental sensitivities. In Sec. 5 we describe implications of our results. We conclude in Sec. 6.
2 Theoretical Model
2.1 Effective interactions
@@
where $m_{\phi}$ is the field’s mass and $\lambda$ is the self-interaction coupling, and the []delimited-[][$\dots$][ … ] indicate possible terms with higher powers of $\phi$. Here, we consider $\phi$ to constitute ULDM with mass $m_{\phi}\lesssim{\rm eV}$ and treat $|\lambda|\ll 1$ as a free parameter that depends on the underlying theory. We focus on parameter space $\lambda>0$, denoting attractive self-interactions between $\phi$ particles [^1] This is also motivated by some theoretical considerations, see e.g. Fan:2016rda .. In Sec. 2.3 we discuss various constraints on $\lambda$.
The couplings of $\phi$ with the SM, and as a consequence their optimal detection methods, will depend on the symmetries such as the parity of the model. Axions and axion-like particles more generally (see e.g. adams2023axion ) are typically parity-odd, which implies derivative couplings to SM fields. On the other hand, parity-even fields (see e.g. Antypas:2022asj ) have different interactions with SM fields. As we discuss below, this leads to distinct detection prospects and promising search strategies.
@@
where $M_{\text{pl}}=1.2\times 10^{19}\,{\rm GeV}$ is the Planck mass and ${\mathcal{O}}_{\rm SM}^{i}$ are dimension-four SM operators labeled by index $i$. In this parametrization the $d_{i}^{(n)}$ couplings are dimensionless. Throughout, we will focus on scalars with quadratic $\propto\phi^{2}$ interactions with SM constituents, corresponding to Eq. (2) with $n=2$. In particular, we consider
$\mathcal{L}_{\text{int}}\supset$ $~{}\frac{4\pi\,\phi^{2}}{M_{\text{pl}}^{2}}\bigg{(}d_{m_{e}}^{(2)}m_{e}\overline{e}e-\frac{d_{e}^{(2)}}{4}F_{\mu\nu}F^{\mu\nu}-\frac{d_{g}^{(2)}\beta(g_{s})}{2g_{s}}G_{\mu\nu}^{a}G^{a\mu\nu}\bigg{)}~{},$ (3)
@@
In the presence of the couplings described in Eq. (3), classical oscillations of the ULDM field at frequency of order $m_{\phi}$ lead to oscillations of the fundamental constants, $m_{e}/m_{p}$ (electron-to-proton mass ratio), $\alpha$ (fine structure constant), and $m_{q}/\Lambda_{\rm QCD}$ (ratio of quark mass to QCD scale), respectively. As we discuss, this allows for unique opportunities to search for their manifestations in quantum sensing experiments. Typically, such searches focus on detecting the local DM abundance in the vicinity of Earth, which is expected to have an energy density of order $\rho_{\rm DM}\simeq 0.4\,{\rm GeV/cm}^{3}$ Read:2014qva . However, burst emissions of relativistic ULDM arising from Galactic transient sources establish novel and unique complementary observational targets Eby:2021ece ; Arakawa:2023gyq . Previously Arakawa:2023gyq , we analyzed the detection prospects for ULDM with linear $\propto\phi$ interactions with the SM constituents, corresponding to $n=1$ case in Eq. (2). In this study, we analyze the signatures and detection prospects for ULDM scalars with quadratic SM couplings.
2.2 Screening
@@
2.3 Constraints on self-interaction coupling $\lambda$
2.3.1 Observational constraints
@@ 2.3.2 Naturalness
Figure 2: One-loop diagrams that generate corrections to ULDM self-interaction coupling $\lambda$ originating from the quadratic couplings to SM $d_{m_{e}}^{(2)}$ (top left), $d_{e}^{(2)}$ (top right) and $d_{g}^{(2)}$ (bottom). See Eq. (3) of text.
@@
$\lambda(\Lambda_{\text{UV}})\simeq$ $~{}\lambda(\mu_{0})-\frac{(d^{(2)}_{m_{e}})^{2}m_{e}^{2}}{4\pi^{2}M_{\text{pl}}^{4}}\Lambda_{\text{UV}}^{2}+\frac{(d^{(2)}_{e})^{2}}{8\pi^{2}M_{\text{pl}}^{4}}\Lambda_{\text{UV}}^{4}+\frac{2(d^{(2)}_{g}\beta(g_{s}))^{2}}{\pi^{2}g_{s}^{2}M_{\text{pl}}^{4}}\Lambda_{\text{UV}}^{4}\,,$ (13)
where $\mu_{0}\ll\Lambda_{\rm UV}$ is a reference energy scale. The same couplings in Eq. (3) imply divergent contributions $\delta m_{\phi}^{2}$ to the the ULDM mass $m_{\phi}$ (see e.g. Craig:2022eqo for a recent review). Requiring a “natural” ULDM mass, with corrections being subdominant $\delta m_{\phi}\lesssim m_{\phi}$, implies
@@
Applying these limits to Eq. (13), the naturalness limits on renormalized self-couplings are
@@
Figure 3: Constraints on ULDM self-interaction $\lambda$ versus mass $m_{\phi}$. The bounds from the Bullet Cluster (cyan) Markevitch:2003at , structure formation (orange) Cembranos:2018ulm , Lyman-$\alpha$ Irsic:2017yje ; Rogers:2020ltq and ultra-faint dwarf (UFD) galaxies Dalal:2022rmp (red vertical thick and dashed lines, respectively), black hole superradiance Baryakhtar:2020gao ; Unal:2020jiy (BHSR, brown), as well as the lower bounds for naturalness arising from $d^{(2)}_{m_{e},e}$ (red), and $d^{(2)}_{g}$ (pink) are shown. The naturalness lines assume benchmark values for the couplings that are allowed by current experimental observations, as labeled. We also illustrate the region where critical boson stars would not undergo bosenova due to general relativistic effects (purple hatched region). The dashed black line represents the benchmark $\lambda_{b}$ that we choose at each mass, as given in Eq. (12) and discussed in the Sec. 2.3. The dotted black line shows the benchmark $\lambda_{b}^{\prime}$ we choose when we consider the BHSR bounds, given in Eq. (30). See text for details.
The constraints from naturalness are shown in Fig. 3 for a benchmark choice of the dilatonic coupling $d_{i}^{(2)}=10^{-4}$, which is allowed by all current experimental bounds. We note that dilatonic couplings that are larger than this are also allowed by experimental constraints at the higher end of the mass spectrum (see Sec. 5 for details), and also the dilatonic couplings could be significantly smaller. The naturalness constraint depends on the choice of dilatonic coupling linearly, as can be seen from Eqs. (17-19). While we illustrate constraints from naturalness considerations, we do not consider them here as strict limits for our choice of benchmark parameter values of self-interaction coupling $\lambda$.
3 Boson Stars and Bosenovae